### v2

Abstract
The increasing frequency of natural and human-induced disasters necessitates advanced visual recognition techniques capable of analyzing critical photographic data. With progress in artificial intelligence and resilient computational systems, rapid and accurate disaster classification has become crucial for efficient rescue operations. However, visual recognition in disaster contexts faces significant challenges due to limited and diverse data from the difficulties in collecting and curating comprehensive, high-quality disaster imagery. Few-Shot Learning (FSL) provides a promising approach to data scarcity, yet current FSL research mainly relies on generic benchmark datasets lacking remote-sensing disaster imagery, limiting its practical effectiveness. Moreover, disaster images exhibit high intra-class variation and inter-class similarity, hindering the performance of conventional metric-based FSL methods. To address these issues, this paper introduces the Attention-based Bhattacharyya-Hellinger Feature Aggregation Network (ATTBHFA-Net), which linearly combines the Bhattacharyya coefficient and Hellinger distances to compare and aggregate feature probability distributions for robust prototype formation. The Bhattacharyya coefficient serves as a contrastive margin that enhances inter-class separability, while the Hellinger distance regularizes same-class alignment. This framework parallels contrastive learning but operates over probability distributions rather than embedded feature points. Furthermore, a Bhattacharyya-Hellinger distance-based contrastive loss is proposed as a distributional counterpart to cosine similarity loss, used jointly with categorical cross-entropy to significantly improve FSL performance. Experiments on four FSL benchmarks and two aerial disaster image datasets demonstrate the superior effectiveness and generalization of ATTBHFA-Net compared to existing approaches. Code repository will be available at https://github.com/GreedYLearner1146/ABHFA-Net.
keywords:
Attention Mechanism, Bhattacharyya Distance, Contrastive Learning, Few-Shot Learning, Hellinger Distance, Image Classification, Variational Learning
1 Introduction
1.1 Computer Vision for Disaster Classification
The epoch of deep learning and Computer Vision (CV) has ushered in transformative advancements, augmenting the effectiveness and efficiency of conventional visual tasks such as scene classification, segmentation, and object detection, particularly within the realm of remote sensing [1]. Although the primary source of remote sensing image data emanates from satellite systems, such as very High Resolution (VHR), Hyperspectral, and Synthetic Aperture Radar (SAR)-based images encompassing an extensive region of interest (RoI) ([2], [3]), recent developments have witnessed the integration of imagery from Unmanned Aerial Vehicles (UAV) and low-altitude aircraft into this image domain. Consequently, these data have become critical in the identification of disasters and subsequent damage assessment. This heightened relevance can be attributed to the escalating trends of global warming [4], which have catalyzed an increase in catastrophic events such as fire outbreaks, floods, landslides, and incidents leading to human injuries or diseases [5]. The diversity and magnitude of such disasters are aptly illustrated in Fig.1 and Fig.2. This refined understanding, buttressed by a synthesis of empirical data and analytical methodologies, promises to enhance our capacities for disaster prediction, response, and mitigation, demonstrating the indubitable value of advancements in deep learning and CV within the context of remote sensing and disaster management.
Figure 1: Some example of image sample from each respective disaster class in the CDD dataset. All the images are adapted from the CDD dataset [6]. Note that there are four non-disaster classes (normal) that are comprised of street images, forest images, sea images, and human (single or group) images.
Figure 2: Some example of image sample from each respective disaster class in the AIDER dataset. All the images are adapted from the AIDER dataset [7]. Note that there is a non-disater (normal) class.
In the realm of deep learning, the efficacy of extant algorithms predominantly hinges upon access to extensive data [8]. A compelling correlation has been established between the depth of the learning model and its performance on the respective visual task, thereby escalating the algorithmic demands during the inference phase. Such requirements present significant challenges for mobile platforms such as Unmanned Aerial Vehicles (UAVs), which operate on limited computational and power resources and necessitate expeditious scene understanding in disaster-stricken areas [9]. Even in scenarios endowed with adequate computational and power resources, data scarcity may compromise the effective implementation of state-of-the-art models during training and inference phases. Consequently, the exigent challenge lies in devising accurate yet efficient classification methodologies that can sustain effectiveness amidst potential data limitations.
1.2 Few-Shot Learning
Few-Shot Learning (FSL) approaches have addressed these concerns by harnessing meta-learning [10, 11, 12], wherein a given model acquires the ability to learn classification through exposure to limited example images from a wide variety of classes. Notably, numerous FSL approaches for CV (for instance, [12], [13], and [14]), employed Convolutional Neural Networks (CNN) for essential feature extraction from input images. Subsequently, these features are processed into an embedding space to devise a distance metric that collocates similar classes to proximate points in the space. Within the sphere of FSL, avant-garde strategies have demonstrated unprecedented classification accuracy in benchmark datasets, superseding traditional non-FSL methods [15]. These strategies distinguish themselves by circumventing exhaustive data processing during training and inference, striking a harmonious equilibrium between efficiency and effectiveness in the face of data constraints. However, it must be noted that the majority of the FSL literature relies on benchmark datasets such as Omniglot [16], miniImageNet [17], tieredImageNet [18], FC-100 [19], and CIFAR-FS [20], which are devoid of disaster-related imagery, thereby curbing the applicability of FSL techniques in pertinent scenarios (although evaluating any novel FSL approaches on the benchmarks is still of upmost importance for establishing a baseline). Nevertheless, the integration of attention-based mechanisms, underpinned by the Matching Network [17], has instigated a surge of innovation in FSL research. Novel contributions such as attention with weight fusion [21], dual attention [22], and self-attention [23], have facilitated the selective processing of salient feature maps while excluding extraneous information. This approach effectively curtails computational costs and enhances model efficiency, thereby laying the groundwork for the advent of refined and potent FSL solutions.
Since a majority of the FSL would often learned from an abundance of base samples with known classes and fine-tuned on unseen new classes, the base classes can be easily recognizable while the new classes may not be predicted accurately if the variance of the query set is high, which is more conspicuous if there are limited query samples or rare classes. This is the high intra-class variance problem and consequently the model may easily confuse the new classes with the base. Simultaneously, there is also the high inter-class similarities issue in which certain class can have common visual features as that of the other classes, such as landslide and collapsed building aftermath, which could exacerbate further confusions.
There are a variety of approaches to reduce inter-class variances: Some instances include data augmentation (e.g., [24], [25]), contrastive learning (e.g.,[26], [27]), and variational learning ([28], [29], [30], [31]). For the latter, the approach involved Variational Autoencoder (VAE) to measure similarities among the embedded feature probability distributions, often utilizing f-divergences such as the Kullback-Leiber (KL) divergence to approximate the latent distribution given the support feature distribution sets. Since the few-shot classification performances are highly dependent on the quality of the support and query set sampling, any sample outlier or anomaly would seriously affect the classification effectiveness. This adds on to the issue as highlighted by Zhang et al. [29] on implementing few-shot models based on feature point estimates, which remained dominant in current FSL literature. To address the former, the Mahalanobis distance can be utilized ([32],[33]), but the metric is more for point-based embedded feature, and assumed that the distributions to be compared had the same covariance matrix. A closely related metric for distribution comparisons is the Bhattacharyya coefficient and distance ([34], [35], [36]), which can be shown to generalize the Mahalanobis distance ([37],[38]), and also has the property of rapid divergence when the class distribution are dissimilar, hence serving as a “push term” to separate dissimilar embedded features distribution during the variational feature aggregation process, analogous to that in contrastive learning but for embedded feature points.
Concurrently, the Hellinger distance [39] is yet another f-divergences which is much less used than the KL divergence. However, in contrast to the Bhattarcharyya coefficient/distance, the Hellinger distance can be considered as a probabilistic analog of the Euclidean distance ([40], [41]), and this means that minimizing the Hellinger distance is equivalent to pulling similar embeddings closer together in a Euclidean-like manner but in the square-root probability space. In a sense this distance serves as a “pull term” in contrast to the Bhattacharyya coefficient as above, once again in analogy to that in contrastive learning. The individual parenthesis term in the Hellinger distance is also bounded between 0 and 1, and this ensures that the value doesn’t diverge to infinity unlike the KL divergence when the denominator probability distribution in the latter goes to zero. There are emerging works on utilizing the Hellinger distance ([42], [43]), but these are applied to the other domain apart from few-shot image classification, until the recent works by Lee et al. ([40], [41]).
1.3 Main Research Gaps and Objectives
However, the role of incorporating attention, which has been more widespread in the point-estimate FSL approaches (and for which some examples have been mentioned previously), has not been focused much upon in variational few-shot image classification (apart from TRIDENT [44]). Consequently, there are little FSL works that explored the combination of attention and the Hellinger (along with Bhattarcharyya coefficient) distance approach for classification in the common dataset benchmark as described, let alone in a remote sensing context (with the exception of [40], [41] in the context of Hellinger distance). As mentioned, this is because most of the variational FSL works focused on its application using the KL divergence. However, we have identified a strong analogous association between the Bhattacharyya coefficient and the Hellinger distance in establishing a contrastive learning paradigm. This means that it is possible to train a variational few-shot model in the same manner in which we trained a contrastive learning classification model, hence enabling us to combine the benefits of both meta-learning approach. Such training can be done via incorporating the respective distance or divergence into the loss function proper, as demonstrated by Wu et al. [45] and Lee et al. [46] in the domain of image dehazing. The latter’s approaches, when modified suitably, can potentially enhance the few-shot capability of the proposed model, making it vital for the aerial disaster classification context, as discussed earlier. Fig. 3 illustrates the analogy between the point feature clustering in ordinary contrastive learning and the Bhattacharyya-Hellinger variational distribution clustering for a three-class scenario.
Figure 3: A pictorial comparison of the point feature-based contrastive learning (left) and the analogous distribution feature-based “contrastive learning” (right) using a three-class scenario, where the classes are denoted by purple circles, maroon diamonds and green hexagons. As mentioned, the Bhattarcharyya coefficient (as a similarity measure) can serve as a “push term” as the distance increases rapidly the more dissimilar the feature distribution, while the Hellinger distance can serve as a “pull-term” since similar embeddings are clustered closer together in a Euclidean-like manner but in the square-root probability space.
Therefore, motivated by the increase in the incorporation of attention mechanism, the possibility of exploring alternative probability distribution-based distance approach for FSL prototype clustering, and the pressing needs of highly adaptable disaster classification scheme, we proposed the Attention-based linearly combined Bhattacharyya and Hellinger Distance-based Feature Aggregation Network (ATTBHFA-Net) for our few-shot feature aggregating paradigm. Our model provides an alternative but novel strategy in feature aggregation for few-shot classification, which retains the encoder architecture (as in the work by [29] and VFA-Net) but rendered the task more effective thanks to the incorporation of attention and a combination of the Bhattacharyya coefficient and Hellinger distance which could serve as a contrastive “push” and “pull” term in the probability distribution space respectively. Since an encoder incorporates bottleneck latent variables $z$, the respective distances between the probability distribution of the feature latent variables $z_{features}$ and the prototype latent variables $z_{prototype}$ are computed. To the best of our knowledge, this is the first time we applied two alternative f-divergence distance measures to the probability distribution of the latent variables and demonstrated such feasibility, all in the context of FSL. In summary, our contributions in this paper are as follows:
  • We proposed ATTBHFA-Net, which not only introduced spatial-channel attention in our few-shot feature embedding procedure, but also incorporated a combined Bhattacharyya-Hellinger distance which serves as a novel VAE strategy in modelling and aggregating the class prototypes in a contrastive learning like-manner. To our knowledge, ATTBHFA-Net is the first of its kind in the domain of FSL which combines attention and the Bhattacharyya-Hellinger distance to perform VAE-based classification on the benchmarks, as well as on remote sensing-based visual disaster classification.
  • To ensure that the variational optimization of our ATTBHFA-Net can be performed in an analogous footing to that of the other related methods (e.g., VAE [47], WVAE [31]), we derived a mathematical relation linking our modified ELBO (which would be known as the $ELBO^{*}$) and the Bhattacharyya coefficient. The corresponding relationship between the modified ELBO and the Hellinger distance was made in [40]. These relations are and have been shown to be mathematically consistent like that of the relations obtained from other variational-based works.
  • We also introduced a Bhattacharyya-Hellinger distance-based contrastive training loss, also known as the Bhattacharyya-Hellinger Softmax loss function $\ell_{BHS}$, which is inspired by the cosine similarity softmax loss in simCLR [48], but with the cosine similarity in the softmax function replaced by our linearly added Bhattacharyya coefficient and Hellinger distance. The overall training loss for our network, $\ell_{ATTBHFA}$, is a weighted linear combination of $\ell_{BHS}$, the ordinary categorical cross entropy loss $\ell_{CCE}$, and the image reconstruction loss $\ell_{rec}$. Such loss combination has been shown to enhance our classification performances relative to the State-Of-The-Arts (SOTA) FSL approaches.
  • We implemented and evaluated selected FSL-based approaches on the respective benchmark datasets (CIFAR-FS [20], FC-100 [49], miniImageNet [17], tieredImageNet [18]), as well as on two disaster-based image dataset, namely, the Aerial Image Database for Emergency Response (AIDER) [7] and the Comprehensive Disaster Dataset (CDD) [6]. The results obtained from the respective dataset in the latter could serve as a baseline in this context of FSL-based classification. The extensive experiments performed on the six datasets demonstrated not only the feasibility of our method but also the ability to generally surpass the other FSLs in terms of the Average Accuracy (AA) classification performance metric. The reconstructed images using the Bhattarcharyya distance component of our model is also shown to outperform that of the VAEs using the other f-divergences.
2 Related Works
In this section, we highlight some of the relevant and existing works in the domain of FSL, before discussing some of its applications to the remote sensing domain.
2.1 On Few-Shot Learning Review
Most of the FSL-based approaches discussed in the following can be categorized into metric-based learning approaches, in which a distanceEV function over the data samples is learned [50]. The objective is to learn an embedding function in the embedded space given a distance function, with each function characterized by the relevant parameters, which is computed by the neural network. The work by Bateni et al. [33] highlighted the importance of selecting the distance metric, which has been largely unfocused on previously due to the common perceptions that non-linear mappings via neural networks should presumably adapt to any kind of metric, and hence making any selected distance choices inconsequential.
One of the first metric-based FSL models manifested in the form of a Siamese network [12]. Here, a pair of input images and its corresponding labels are selected at random from a dataset, and the similarity between the images and labels is learned in the training phase. A similarity function and a loss function are commonly employed, and for a Siamese network, the contrastive loss function ($\ell_{contrastive}$) is utilized in training the model to identify similarities. The prototypical network (ProtoNet), introduced by [14], differs from the Siamese network in that it processes the image features into a higher-dimensional embedding space instead of directly from the images, for which samples from the same class can be easily clustered together. For each class cluster, a class prototype is calculated as the average of the data points in that cluster. Relation Network (RelationNet) [51] computes relation scores, which compare the feature maps of the query sample set and the support set, and has the advantage unlike that of the Siamese and ProtoNet in that the similarity metric needs not be defined manually, but is determined from the CNN feature extractor directly. Also, RelationNet is unique relative to the other FSL approaches in that the Mean-Squared Error loss ($\ell_{MSE}$) was used instead of the cross-entropy loss. This is because the output, as mentioned, is a relation score that warrants the problem to be modeled as a regression problem instead. The Matching Network (MatchingNet) [17] utilizes the attention kernel and cosine similarity in embedding of the support and query examples, and removes the need of fine-tuning for adaptation to new classes.
Transductive Information Maximization (TIM) was introduced by [52] which fine-tunes the class prototypes based on the classification error in the support images, as well as any Mutual Information (MI) between the query features and the label predictions. Euclidean distances are used to guide the classification and its subsequent updates. TIM, as the name suggests, maximizes the MI between query features and the label predictions for a given few-shot task at the inference phase, while minimizing the cross-entropy loss on the support set. The method is transductive in that the model classifies any unlabeled query examples for a given few-shot training task all at once, instead of classifying the sample one at a time, as in the usual inductive methods. A transductive fine-tuning approach was proposed by [53] which performs fine-tuning of the training parameters based on the classification error incurred in the support images and the classification entropy on the query images. Like TIM, Euclidean distances are utilized to guide the classification of the prototypes. Another fine-tuning approach, which uses class prototypes, was proposed by [10] and like transductive fine-tuning, the classification error incurred on the support images is used in the fine-tuning process. The query prototypes are also classified and updated based on cosine distances, similar to [54].
Figure 4: Pictorial depiction of two support feature encoding approaches for a three-class ($c_{1},c_{2},c_{3}$) clustering setting. The grey dot represents the centroid (mean) of each of the class prototype ($\mu_{1},\mu_{2},\mu_{3}$), with its sample points denoted by red, green and blue respectively. The “ellipses” on the right represents the probability distributions of the samples from each distribution. The pale-red dot in the center represents a feature sample point, with the dotted lines denoting its distances to the centroids.
All of the metric-based FSL methods described previously make use of the query and support feature point in the embedded space directly for the computation and clustering. As highlighted by Zhang et al. [29], there are two inherent challenges with such approaches. The first is that estimating a particular feature point in the embedded space can be rendered difficult when the support feature points are limited and unevenly distributed in the space. The second is the lack of interpretability as a single embedding does not suffice to indicate a class. The first point implies that computing the mean of the support features from each class can be heavily influenced by external noise, and the second point implies that utilizing distributions instead of points gives better representational and descriptive power. One solution is to utilize variational distribution estimation, as Zhang et al. [29], and VFA-Net set out to do. The differences between a feature point-based and distribution-based encoding methodology are depicted in Fig.4. Such a network, as mentioned in the introductory section, is inspired by the need to reduce the inherent variances presented in the dataset and increase the model’s robustness to such variances. By modeling each class’s common distribution via a VAE [55], followed by training on a huge amount of base classes and fine-tuning on the novel examples, the learned inter-class variances can be transferred to the new classes, and thus achieving effective estimation of the new classes’ distributions using only limited examples. Robust predictions are then implemented via sampling support features and aggregating them with the query features. The Bhattacharyya and the Hellinger distance in our ATTBHFA-Net shares similar characteristics with the KL divergences in [29] and VFA-Net (as well as the Wasserstein distance in [31]) in that sample feature point distribution in the bottleneck layer is compared to a referenced probability distribution, and that both the mean values and standard deviation of the respective sampled features and prototypes are required. However, the VAE-KL divergences does not effectively account for possible outlier-based data features (which is the inter-class similarity problem that manifests in the form of supposedly non-disaster (normal) images displaying some level of disaster-based features) during the few-shot sampling process. Therefore, utilizing our ATTBHFA-Net’s Bhattarcharyya-Hellinger distance approach can better group these data into the respective disaster classes with higher probability, thanks to the Hellinger distance’s capability to take into account feature outliers, hence reducing mis-classification. Together with the capability to model latent representation distribution in analogous with that of the Euclidean distance for point-estimate features (since the Bhattarcharyya-Hellinger distance aggregates the feature distribution in a contrastive learning-like manner), the aforementioned traits motivated us to incorporate the combined distance for feature aggregation in our work.
2.2 On Applications to Remote Sensing
Few-shot learning has been applied in the area of classification, semantic segmentation, as well as object detection in remote sensing. Some of the datasets utilized include NWPU-RESISC45, WHU-RS19, UC Merced, AID, and OPTIMAL-31 [56]. Some systematic reviews of few-shot learning in remote sensing has been put forward by Lee et al. [3], Qiu et al. [57] and Li et al. [58], but here we will briefly review a few relevant works.
Lang et al. [59] introduced a global rectification and decoupled registration for satellite few-shot image segmentation that made use of several globally updated dynamic prototypes to provide extra non-target segmentation cues, as well as to avoid erroneous prototype activation caused by query-support image pairings’ unpredictability. In order to attain more focused and customized item localization, the foreground and background information flows are separated, preventing needless confusion caused by redundant information. Cheng et al. [60] introduced a holistic prototype activation network to address the issue of overfitting in base classes that produces inaccurate segmentation boundaries in satellite imagery. It comprised of a training-free scheme to extract the prior category representations, a prototype activation module that produces dependable activation maps and well-matched query characteristics through redundant object and class filtering, and a Cross-Referenced Decoder for multi-level feature aggregation and interactive feature re-weighting. The authors verified the model’s versatility via evaluation on dataset with weak-label, zero-shot segmentation, and video object segmentation.
Task-Adaptive Embedding Network (TAE-Net) was proposed by Huang et al.[61] for few-shot satellite image classification, in which a feature encoder learn the embedded features for pre-training. An adaptive task-adaptive attention module is used to select the essential embedded characteristics during learning, and the meta-trained network predict the query set. Kim et al. [62] introduces the Self-Attention Feature Selection Network (SAFFNet) which utilized a self-attention module to infuse features across several scales, similar to a spatial pyramid network. To improve the match between features from the query set and the combined features in the support set, the Self-Attention Feature Selection (SAFS) module is used. Yuan et al. [63] presented a Multi-Attention Deep Earth Mover Distance approach that simultaneously attend to the global and local features for classification. The local attention module captures important and subtle local elements while suppressing others, which improves performance and lessens tiny and big intra-class inequalities.
All of the aforementioned works utilized point-based features instead of distributive-based features. Recently, Wang et al. [64] proposed a variational meta-learning paradigm for remote sensing image classification that enables the model to adapt to different tasks via variational inference and parameter generation via the “tasks central points”, with the “central points” obtained from the distributions of meta-parameters within the paradigm. However, the method is not evaluated in a few-shot context, and there seems to be a lack of variational few-shot classification work on the RGB spectrum in remote sensing, although there exists a related work using the hyperspectral spectrum [65]. Hence, the proposed method as follows will serve as a pioneering work on incorporating alternate probability distribution divergence/distance measures for variational few-shot classification, and applies it in the remote sensing scenario, specifically for disaster classification from an aerial perspective via UAVs.
3 Formal Problem Definition
3.1 Few-Shot Formal Problem Definition
For the few-shot classification exposition, we posit the existence of a labeled training set designated as $X_{base}=\{x_{i},y_{i}\}_{i=1}^{N_{base}}$. Here, the symbol $x_{i}$ represents the raw features pertaining to the $i^{th}$ sample, while $y_{i}$ signifies the one-hot encoded label associated with the same sample. Within the scope of FSL literature, this labeled ensemble is known as the meta-training (or base) dataset.
Adhering to the FSL paradigm, we consider a test dataset $X_{test}=\{x_{i},y_{i}\}_{i=1}^{N_{test}}$ , comprising a novel set of classes represented by $Y_{test}$. The uniqueness of this set is encapsulated in the equation $Y_{base}\cap Y_{test}=\emptyset$, illustrating that $Y_{test}$ introduces completely new class entities not included in $Y_{base}$. The few-shot scenario involves the creation of randomly sampled few-shot tasks from this test dataset, each encompassing a limited number of labeled instances. In this context, each $k$-way-NN-shot task encompasses the random selection of $N_{s}$ labeled instances from each of the $k$ different classes. The set of these labeled instances, represented by $S$, forms the support set, characterized by its size as $|S|=N_{s}\cdot K$. Alongside this, each task also possesses a query set denoted as $Q$, consisting of $|Q|=N_{Q}\cdot K$ unlabeled (or novel) instances from each of the $K$ classes. The training is reiterated with different image batches presented to the model, and once it is done on the base set, the labeled support sets are utilized for task-specific adaptation. The effectiveness of these techniques is then assessed based on their performance on the unlabeled query sets, providing a measure of the model’s capacity to generalize to new data categories based on limited examples.
3.2 Variational Inference Problem Definition
In VAE, the goal is to estimate the true distribution $p(\bf{x})$ of the data distribution, which is not known a priori. However, we can utilized Bayes’ rule to determine $p(\bf{x})$ via determining the latent distribution $p(\bf{z})$, as well as the conditional distributions associated with the encoder and the decoder $p(\bf{z}|\bf{x})$ and $p(\bf{x}|\bf{z})$, respectively. Although $p(\bf{z})$ can be approximated via a unit variance Gaussian $\mathcal{N}(0,1)$, $p(\bf{z}|\bf{x})$ and $p(\bf{x}|\bf{z})$ are also unknown variables, which can be approximated via imposing proxy distributions $q_{\phi}(\bf{z}|\bf{x)}$ (for the encoder) and $p_{\theta}(\bf{x}|\bf{z})$ (for the decoder), where ($\theta$$\phi$) are the optimizing variables to be determined via training. The corresponding objective function to be determined in this aspect is the Evidence Lower Bound (ELBO) which serves as lower bound for the prior (true) distribution $p(\bf{x})$ via
$\textrm{log}(p(x))\geq ELBO=\int\bf{q_{\phi}(\bf{z}|\bf{x})}\textrm{log}\left(\frac{p_{\theta}(\bf{x},\bf{z})}{q_{\phi}(\bf{z}|\bf{x})}\right)dz.$ (1)
By maximizing the ELBO, we can attain the goal of maximizing the logarithm of $p(\bf{x})$. The equality becomes equal when $q_{\phi}(\bf{z}|\bf{x})$ matches the true distribution of $p(\bf{z}|\bf{x})$. The relationship between the ELBO and the KL divergence is very well established ([29], [66], [47]), and the same can be said for the corresponding relationship with the Wasserstein distance [31]. The goal of this work is to establish a mathematically valid relationship between the modified ELBOs and the Bhattarcharyya coefficient along with the Hellinger distance.
4 Our Approaches
In this section, we elaborate on the various key aspects of our ATTBHFA-Net, as well as the experimental setup and procedures for the evaluation of the FSL SOTAs relative to our proposed design. The loss functions and evaluation metrics utilized throughout our experiments are also highlighted.
4.1 Components of ATTBHFA-Net
There are four novel key components in our ATTBHFA-Net: the incorporation of channel-spatial attention in the encoder, the Bhattarcharyya coefficient measure, the Hellinger distance measure, and the Bhattarcharyya-Hellinger Softmax ($\ell_{BHS}$) loss function.
4.1.1 Attention Mechanism
Figure 5: An illustration of the attention model incorporated in the encoder and decoder component of our ATTBHFA-Net. The architecture comprised of the channel attention and the spatial attention mechanism in which the features are processed consecutively. Note the skip connection between the original and the respective refined features in both the channel and the spatial attention mechanism.
The attention module we have incorporated into the encoder architecture is composed of channel attention and spatial attention modules. For channel attention, the spatial dimensions of the intermediate feature map $\boldsymbol{\phi}$ are extracted and squeezed using a combination of average and max pooling. The output features from the respective pooling processes are inputted into a Multi-Layer Perceptron (MLP) network, and they are finally combined via an element-wise summation. The aforementioned processes is mathematically described as
$\begin{split}\boldsymbol{M_{c}(\boldsymbol{\phi})}&=\sigma(MLP(AvgPool(\boldsymbol{\phi})+MLP(MaxPool(\boldsymbol{\phi})))\\ &=\sigma(\boldsymbol{\omega_{1}}(\boldsymbol{\omega_{0}}(\boldsymbol{{\phi^{c}}_{avg}}))+\boldsymbol{\omega_{1}}(\boldsymbol{\omega_{0}}(\boldsymbol{{\phi^{c}}_{max}}))),\end{split}$ (2)
where $\sigma$ is the sigmoid (activation) function, $\boldsymbol{\omega_{0}}$, $\boldsymbol{\omega_{1}}$ are the MLPs weights, $\boldsymbol{M_{c}(\boldsymbol{\phi})}$ is the channel attention map, and $\boldsymbol{{\phi^{c}}_{avg}}$ and $\boldsymbol{{\phi^{c}}_{max}}$ are the feature maps obtained using the aforementioned pooling operations.
For spatial attention, both aforementioned pooling operations are implemented again, followed by applying a convolutional layer of a filter size of 7 along with the sigmoid activation function to produce the spatial attention map. This is quantitatively described as
$\begin{split}\boldsymbol{M_{s}(\phi)}&=\sigma(f^{7\times 7}([AvgPool(\boldsymbol{\phi});MaxPool(\boldsymbol{\phi})]))\\ &=\sigma(f^{7\times 7}([\boldsymbol{{\phi^{s}}_{avg}};\boldsymbol{{\phi^{s}}_{max}}])),\end{split}$ (3)
where $\boldsymbol{M_{s}(\phi)}$ denotes the spatial attention map, and $\boldsymbol{{\phi^{s}}_{avg}}$ and $\boldsymbol{{\phi^{s}}_{max}}$ denote the feature maps obtained using the aforementioned pooling operations respectively.
Altogether, the spatial and channel attention maps are processed as
[…]
where $\otimes$ is the tensor product between the extracted features ($\boldsymbol{\phi^{\prime}}$ and $\boldsymbol{\phi^{\prime\prime}}$) and the attention maps. The overall architecture of our attention mechanism is illustrated in Fig.5.
In the next two subsection we make some subtle notational changes to formulate the VAE framework into the few-shot learning context. The proxy distributions $q_{\phi}(\bf{z}|\bf{x})$ and $p_{\theta}(\bf{z}|\bf{x})$ now become $q_{\phi}(\bf{z}|\mathcal{S})$ and $p_{\theta}(\bf{z}|\mathcal{T},\mathcal{S})$, where $\mathcal{T}$, $\mathcal{S}$ represents the training distribution and the support set respectively. For the latter distribution we also write it as $p_{\theta}(\bf{z}|\mathcal{T})$ and implicitly implied that $\mathcal{S}$ is a subset of $\mathcal{T}$.
4.1.2 Bhattarcharyya Distance
The Bhattarcharyya distance $D_{B}$ in our ATTBHFA-Net is mathematically depicted as
$D_{B}=-\textrm{ln}\left(\int\sqrt{(p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz\right).$ (5)
The integrand with the product term $\int\sqrt{(p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S}))}dz$ is also known as the Bhattarcharyya Coefficient ($BC$). We make the assumption that $p_{\theta}(z|\mathcal{T})$ manifested as a multivariate Gaussian $\mathcal{N}(\mu_{i},\sigma_{i})$, whereby the mean is $\mu_{i}$ and the standard deviation is $\sigma_{i}$, and the (proxy) posterior distribution $q_{\phi}(z|\mathcal{S})$ also took on a similar distribution but with a mean of $\mu_{j}$ and a standard deviation of $\sigma_{j}$ (i.e., $\mathcal{N}(\mu_{j},\sigma_{j})$). The value of the $\mu$ and $\sigma$ of the support feature set can be extracted from our attention-based encoder network $\mathcal{F}_{encoder}(S)$ after extracting its latent representation $\boldsymbol{z}$ and calculating the distributions.
To ensure that the variation optimization can be done on a similar footing as that of the KL divergence and Wasserstein distance, we can derived a modified Evidence Lower Bound ($ELBO^{*}$) which is related to the Bhattarcharyya distance (and hence coefficient) in a similar manner as that of Lee et al. [40]. The mathematical derivation is detailed in the appendix section, but we denote the final mathematical result as follows:
$\begin{split}ELBO^{*}=\int q_{\phi}(z|\mathcal{S})\textrm{ln}\left(p_{\theta}(z|\mathcal{T})\right)dz+2D_{B}\\ =\int q_{\phi}(z|\mathcal{S})\textrm{ln}\left(p_{\theta}(z|\mathcal{T})\right)dz-2ln(BC)\end{split}$ (6)
where $\int q_{\phi}(z|\mathcal{S})\textrm{ln}\left(p_{\theta}(\mathcal{T}|z)\right)dz$ is the reconstruction term and $ELBO^{*}$ can be expressed as $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-ELBO$. It can be shown that if $p_{\theta}(z|\mathcal{T})=q_{\phi}(z|\mathcal{S})$, $ELBO^{*}$ is equivalent to the reconstruction term, in agreement with previous related work that associates the respective ELBO with the KL divergence ([29], [66]), the Wasserstein distance [31] and the Hellinger distance [40].
4.1.3 Hellinger Distance
The (square of the) Hellinger distance $D_{H}$ measure in our ATTBHFA-Net is mathematically depicted as
$D_{H}^{2}=\frac{1}{2}\int\left(\sqrt{p_{\theta}(z|\mathcal{T})}-\sqrt{q_{\phi}(z|\mathcal{S})}\right)^{2}dz.$ (7)
The Hellinger distance can better handle outliers. Unlike the KL divergence ($D_{KL}$) which can amplifies the small errors associated with the outliers in the low-probability region (occurring at both ends of the distribution tails and hence contributes to larger penalties), the Hellinger distance diminishes the effect of the large differences in the error and thus contributing to smaller penalties. Also, unlike $D_{KL}$, $D_{H}$ will not diverge when any input probability distribution tend very close to zero, thus the problem of $D_{KL}$ value diverging to infinity is not a concern in $D_{H}$. Finally, the Bhattarcharyya coefficient can be linked with the Hellinger distance via $D_{H}^{2}=1-BC$, and from this we can clearly see the opposing relations between the two metric in the analogous contrastive learning of our model: Since $BC$ increases as the feature distribution becomes more dissimilar, the $D_{H}^{2}$ decreases as the seemingly dissimilar features would eventually “meet” their similar counterparts in the embedded space, hence achieving the aggregation effect.
As derived from [40] and [41], $D_{H}^{2}$ can be related to the modified ELBO as
$\begin{split}ELBO^{\prime}=\int\textrm{log}(p_{\theta}(\mathcal{T}))dz+log\left(1-D^{2}_{H}\right)^{2},\end{split}$ (8)
where $ELBO^{\prime}$ is the new ELBO term expressed as $\frac{ELBO}{q_{\phi}(z|\mathcal{S})}$ in the original paper, and $\int log(p_{\theta}(\mathcal{T}))dz$ is the evidence term. However, in the appendix we will derive the aforementioned relationship again from another mathematical angle, and arrived at
$\begin{split}ELBO^{\prime\prime}=-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+log(1-D_{H}^{2})^{2}\end{split}$ (9)
where $ELBO^{\prime\prime}=ELBO-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$. We note that
$\begin{split}\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz=\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T})p_{\theta}(\mathcal{T}))dz=\\ \int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+\int q_{\phi}(z|\mathcal{S})p_{\theta}(\mathcal{T})dz.\end{split}$ (10)
So we can rewrite $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz$ as $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz+\int q_{\phi}(z|\mathcal{S})p_{\theta}(\mathcal{T})dz$. Move the term $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ to the left-hand side of equation 9 and the term $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ cancels out in $ELBO^{\prime\prime}$. By dividing both sides by $q_{\phi}(z|\mathcal{S})$, we recovered equation 8.
Together, we can express the optimization objective of the ATTBHFA-Net (to maximize $ELBO^{*}$ and $ELBO^{\prime\prime}$ (equivalent to minimizing the Bhattarcharyya coefficient and Hellinger distance) as
$\underset{\phi,\theta}{\arg\max}\sum_{x\in\mathcal{\mathcal{T},\mathcal{S}}}ELBO^{*}+\underset{\phi,\theta}{\arg\max}\sum_{x\in\mathcal{T},\mathcal{S}}ELBO^{\prime\prime}.$ (11)
4.1.4 Bhattarcharyya-Hellinger Softmax (BHS) Loss
Our proposed new training loss function also consisted of BHS loss $\ell_{BHS}$. As mentioned in the introductory section, the modification made here is that instead of the cosine similarity, the Bhattarcharyya coefficient parenthesis term along with the Hellinger similarity are combined and directly incorporated into the parentheses of the softmax function as highlighted below:
$\ell_{BHS}=-\textrm{log}\left(\frac{\textrm{exp}((BC^{\prime}+\{v_{(Q,i)},c_{(Q,j)}\})/\tau)}{\sum_{i=1}^{2N}\textrm{exp}((BC^{\prime}+\{v_{(Q,i)},c_{(Q,j)}\})/\tau)}\right),$ (12)
where $\tau$ denotes the temperature parameter of the softmax function, set as 0.01 throughout our training. Here, we denoted $\{v_{(Q,i)},c_{(Q,j)}\}$ to be the Hellinger similarity between the query distribution $i$ and the class prototype distribution $j$ which is mathematically described as $\left(\sqrt{p_{\theta}(z_{i}|\mathcal{T,Q})}-\sqrt{q_{\phi}(z_{j}|\mathcal{T,C})}\right)$, where now the distributions are conditioned on the query set and class prototypes. The $BC^{\prime}$ is defined as the parenthesis term inside $BC$, which once again is conditioned on the query set and class prototype distribution as $\sqrt{p_{\theta}(z_{i}|\mathcal{T,Q})q_{\phi}(z_{j}|\mathcal{T,C})}$.
However, during our simulation, we have found that utilizing the loss function in the mathematical form above would lead to a gradient explosion due to the instability incurred when multiplying two probability distributions together in $BC^{\prime}$, in accordance to its definition. To address this problem, we recalled the relationship between the Bhattarcharyya coefficient and the Hellinger distance as described in the previous section: $BC=1-D_{H}^{2}$, which implies that $\int\sqrt{p_{\theta}(z_{i}|\mathcal{T,Q})q_{\phi}(z_{j}|\mathcal{T,C})}dz=1-\frac{1}{2}\int\left(\sqrt{p_{\theta}(z_{i}|\mathcal{T,Q})}-\sqrt{q_{\phi}(z_{j}|\mathcal{T,C})}\right)^{2}dz$. We can now simply take the derivative of term with respect to $dz$ to obtain $\sqrt{p_{\theta}(z_{i}|\mathcal{T,Q})q_{\phi}(z_{j}|\mathcal{T,C})}=-\frac{1}{2}\left(\sqrt{p_{\theta}(z_{i}|\mathcal{T,Q})}-\sqrt{q_{\phi}(z_{j}|\mathcal{T,C})}\right)^{2}$, as $\frac{d}{dz}(1)=0$. Therefore, the Hellinger similarity is related to $BC^{\prime}$ by the negative-halve of the former. Consequently, equation 12 can be rewritten as
$\ell_{BHS}=-\textrm{log}\left(\frac{\textrm{exp}((-0.5\{v_{(Q,i)},c_{(Q,j)}\}^{2}+\{v_{(Q,i)},c_{(Q,j)}\})/\tau)}{\sum_{i=1}^{2N}\textrm{exp}((-0.5\{v_{(Q,i)},c_{(Q,j)}\}^{2}+\{v_{(Q,i)},c_{(Q,j)}\})/\tau)}\right)$ (13)
$=-\textrm{log}\left(\frac{\textrm{exp}((\{v_{(Q,i)},c_{(Q,j)}\}(1-0.5\{v_{(Q,i)},c_{(Q,j)}\}))/\tau)}{\sum_{i=1}^{2N}\textrm{exp}((\{v_{(Q,i)},c_{(Q,j)}\}(1-0.5\{v_{(Q,i)},c_{(Q,j)}\}))/\tau))}\right),$ (14)
Note that our $\ell_{BHS}$ closely resembled that of the ordinary negative log likelihood loss function $\ell_{nll}$ used for multi-class classification. The motivation of such loss is that it generalizes the softmax loss not only in the simCLR, but also in the Hellinger similarity softmax loss function as laid out in [40] and [41]. As the Bhattarcharyya coefficient and Hellinger distance are complementary in the contrastive learning-based distribution aggregation, the contrastive power of the loss function can be refined further via our novel $\ell_{BHS}$. This can be made conspicuous by the $1-0.5\{v_{(Q,i)},c_{(Q,j)}\}^{2}$ term which implies the penalization of any query pairs that are not similar in the square of the Hellinger distance measurement, and consequently the increment in the Bhattarcharyya coefficient. Similar to the cosine similarity and the Hellinger similarity softmax loss, in the denominator, we sum the numbers in parentheses over the number of positive samples and divided by two times the batch size $N$ selected per training epoch (i.e., $2N$).
Figure 6: The illustration of our ATTBHFA-Net architecture. It receives a set of support and query images as input, which are then processed by an attention-based encoder to extract their features. These features undergo encoding into latent representation in which the respective Gaussian probability distributions are obtained. The Bhattarcharyya coefficient and Hellinger distance computation process involves deriving confidence spaces for each class prototype’s distribution and comparing their overlapping degree with feature point distribution (represented by a red dot within a red dotted circle). The classes’ maximum probabilities are leveraged to generate a confidence score, indicating the predicted label for a given sample. To perform training via the Bhattacharyya-Hellinger Softmax loss $\ell_{BHS}$ and categorical cross-entropy loss $\ell_{CCE}$, we use ATTBHFA-Net loss: $\ell_{ATTBHFA}$. $y_{\mathcal{S}}$ and $y_{\mathcal{Q}}$ denote the ground-truth label from the support and query set image respectively, and $\hat{y}_{S^{\prime}}$, $\hat{y}_{Q^{\prime}}$ denote the predicted label from the support and query set image respectively. The four quantities are relevant for computing the $\ell_{CCE}$. To perform $\ell_{rec}$, the reconstructed support and query image set $\mathcal{S^{\prime}}$, $\mathcal{Q^{\prime}}$ are compared with the original support and query set image $\mathcal{S}$, $\mathcal{Q}$. The diagram presented above illustrates a 5-way-1-shot Fast FSL training method.
4.2 ATTBHFA-Net Architecture
Together, our ATTBHFA-Net architecture is illustrated in Fig.6, which also clearly depicts the aforementioned components. Our ATTBHFA-Net first involved extracting and encoding the input feature with the attention mechanism, $\mathcal{F}_{encoder}$, to the support set $S$ and the query set $Q$ separately. The respective $\mu$ and $\sigma$ are obtained respectively, and the bottleneck features $\boldsymbol{z}$ are directly obtained via deploying the re-parameterization trick [47] $z=\mu+\zeta\odot\sigma$, where $\zeta$ is the normal distribution and $\odot$ is the hadamard product (element-wise multiplication). The Gaussian distribution $\boldsymbol{p}$ and $\boldsymbol{q}$ are then obtained. The relevant Bhattarcharyya coefficient and Hellinger distance computation then proceeds. This comprises the comparison of the selected feature distribution in the embedding space (denoted by a red cross surrounded by the red dotted circle) in Fig.6) with that of the prototypes’ feature points by measuring the degree of overlap between the respective distributions. These series of steps closely mimicked that of the KL divergence, Wasserstein distance, and Hellinger distance-based VAEs, except with the addition of our Bhattarcharyya-Hellinger probability comparison structure. The respective class-specific distributions that are estimated from the sets are utilized to calculate the probability of the target data, which is the maximum probabilistic attainable among the class. This is then used to generate a confidence score for the label to which the sample is most likely to belong (predicted label), in agreement with [29]. Throughout the above processes, the resultant and (reconstructed) support $S^{\prime}$ and the query’s probabilistic features $Q^{\prime}$ are guided by the loss functions $\ell_{BHS}$, $\ell_{CCE}$ and $\ell_{rec}$, which are combined for the network training on the sets. Therefore, the total loss for our model training $\ell_{ATTBHFA}$ is described by
$\ell_{ATTBHFA}=\lambda_{1}\cdot\ell_{BHS}+\lambda_{2}\cdot\ell_{CCE}+\lambda_{3}\cdot\ell_{rec},$ (15)
in which $\ell_{CCE}$ is mathematically described by
$\begin{split}\ell_{CCE}=-\sum_{i}^{Q}y_{i}log\left(p(\hat{y}_{i}=y_{i}|\mathcal{T}_{meta})\right),\end{split}$ (16)
and the reconstructed loss $\ell_{rec}$ being mathematically described by
[…]
where $y_{i}$ and $\hat{y}_{i}$ represent the ground-truth and predicted labels respectively, $\mathcal{T}_{meta}$ denotes a meta-task and runs through the total number of queries $Q$ per training batch in our experiment, $\mathcal{F}_{dec}$ denotes the decoder backbone, $(\cdot)||$ being the L1-norm, $S$ and $S^{\prime}$ being the original and reconstructed support set respectively, and lastly $Q$ and $Q^{\prime}$ being the original and reconstructed query set respectively. The $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are hyperparameters which are set to 0.7, 0.3 and 1.0 respectively for the benchmark dataset training, and 1.0, 0.5, and 1.0 for the disaster image classification dataset.
5 Experimental Procedures
This section is split into three parts: the first part concerns the benchmark datasets utilized for our model evaluation along with the SOTAs, the second part concerns the selected aerial disaster image classification datasets, once again utilized for the model evaluation along with (another set of) SOTAs, and the last part concerns the relevant configurations and hyperparameters.
5.1 On Benchmark Datasets
As mentioned in the introductory section, we evaluated our approach relative to the SOTAs FSL approaches on the FC-100, CIFAR-FS, miniImageNet, and tieredImageNet. The selected SOTAs for each dataset is the same as that laid out in Roy et al.’s work [67], except that we included the KL few-shot VAE by Zhang et al. [29] (We abbreviated it as VFS throughout the numerical table of results) and the Hellinger few-shot VAE by Lee et al. ([40], [41]).
The FC-100 [19] have 60 meta-training classes, 20 meta-validation classes, and 20 meta-testing classes, each with 600 images of size 32 $\times$ 32. Few-shot learning can be made difficult in this dataset as the small image size are correlated with poor image quality.
The CIFAR-FS [49] have 20 meta-testing classes, 16 meta-validation classes, and 64 meta-training classes. Like the FC-100, the images also have a size of 32 $\times$ 32, and the few-shot tasks are once again rendered difficult due to the significant intra-class similarity between the images.
In the miniImageNet dataset [17], there are 100 classes each with 600 images and each with an image size of 84 $\times$ 84. There are 64 meta-training classes, 16 meta-validation classes, and 20 meta-testing classes. Lastly, the tieredImageNet [18] included 351 classes for meta-training, 97 classes for meta-validation, and 160 classes for meta-testing, once again with each image size of 84 $\times$ 84.
5.2 On Aerial Disaster Image Classification Datasets
We have selected the triplet network, ProtoNet, RelationNet, MatchingNet, BD-CSPN, SimpleShot, prototype Fine-Tuning (FT), transductive FT, TIM, and finally the HELA-VFA for evaluation against out ATTBHFA-Net in this experiment. Unlike that of the benchmark dataset, we do not include ANROT-HELANet in this evaluation, as corrupting the input images with adversarial noises is unrealistic and undesirable in the execution of UAV visual search-and-rescue, since such noises are generated artificially. As mentioned in the introductory section, we utilized the AIDER and the CDD dataset for the algorithmic baseline evaluation. For AIDER, the dataset is divided into a train-valid-test ratio of 4:1:2. Except as otherwise indicated, the ResNet12 architecture [68] is used as the CNN backbone throughout our experiment. Due to the limited amount of classes available for an effective meta-training and -testing class split, we performed “supervised few-shot learning” in the sense of teaching the model to recognize limited amount of sample from each class during training and generalizing them to the same limited amount from the testing subset at inference time. As for the CDD, the standard few-shot learning paradigm applies. Also for the latter, we did not apply any under-sampling, despite the obvious class imbalances as depicted in Table 2. This is because one of the class in the meta-test set (earthquake) is of a very small amount relative to the rest. If the majority classes are undersampled, many useful information regarding the other test classes would be significantly lost, leading to underfitting [69].
AIDER [7]
This data set consisted of aerial images belonging to any of the four types of disasters, namely the aftermath of collapsed buildings, fires, floods, and traffic accidents, as well as the non-disaster type (labeled as normal). The latter class contained a significantly higher number of data samples than the disaster classes to replicate the real-world scenario as much as possible. The author only made a subset of the original dataset publicly available, and it contained 511 collapsed building images, 521 fire images, 526 flood images, 485 traffic images. and 4390 normal images, making a total of 6433 images, as illustrated in Table 1. The image sizes provided vary across the data set, but in our experimental approach we cropped them all to $\times128$$\times$3.
Table 1: The training, validation and test image sets for each AIDER subset class.
Class Train Valid Test Total per Class
Collapsed Building 367 41 103 511
Fire 249 63 209 521
Flood 252 63 211 526
Traffic 232 59 194 485
Normal 2107 527 1756 4390
Total Per Set 3207 753 2473 6433
CDD [6]
As shown in the introductory section, the CDD comprised of 7 disaster classes and 4 non-disaster classes. This dataset contained not only the classes found in AIDER, but also new disaster classes such as land disasters (e.g., drought and landslides). For the meta-training dataset, we chose sea, street/undamaged buildings, and human for the non-disaster classes, and landslide, wildfire, and infrastructure damages as the disaster classes. The meta-testing dataset includes water disaster, urban fire, drought, earthquake, and forest (non-disaster) class. Note that at least 5 testing classes are required for 5-way-5-shot evaluation. Table 2 depicts the number of samples in each meta-train and meta-test class. Like AIDER, all images used are cropped to $\times128$$\times$3.
Table 2: The meta-training and testing image sets for the CDD classes.
Meta-Train Class No.of Images Meta-Test Class No.of Images
Infrastructure Damages 1418 Earthquake 36
Landslide 456 Drought 201
Wild Fire 514 Urban Fire 419
(Non-Damage) Sea 2274 Water Disaster 1035
(Non-Damage) Streets and Buildings 4572 (Non-Damage) Forest 2271
(Non-Damage) Human 120 - -
Total Training Images 9354 Total Testing Images 3962
Although some of the CDD images do not look like aerial images at first glance, especially images of the normal and infrastructure damage class, we need to recall the ability of a UAV (rotary-based) to fly and hover at a lower altitude that allows for prolonged surveillance of the region of interest. Therefore, seemingly non-aerial images could be taken by a UAV hovering at low altitude, and such images could also serve as a reference for UAV visual classification at the aforementioned altitude.
5.3 Experimental Configurations and Hyperparameters
All images in the respective dataset are re-scaled via division of each original pixel value by 255 as a pre-processing step. All experimental simulations for this work were performed using PyTorch library in Python utilizing Google’s Colab Pro+ Tesla T4 Graphical Processing Units (GPU). The number of training episodes utilized for our approach is 80000, and the optimizer utilized for our network is that of Adam with a learning rate of 0.001. For the temperature value in $\ell_{BHS}$, we set it to 0.01.
Evaluation Metrics
For all the datasets, we used the average classification accuracy scores as the main metric to assess the classification effectiveness of each SOTAs, as well as our approach. Average Accuracy (AA) is mathematically defined as
$\begin{split}AA=\sum_{i=1}^{N}\frac{TP_{i}+TN_{i}}{TP_{i}+FN_{i}+TN_{i}+FN_{i}},\end{split}$ (19)
where the variables $TP$, FPFP, TNTN, and $FN$ represent the true positive, false positive, true negative, and false negative classes, respectively, and $N$ denote the total number of classes. In addition, we also utilized the confusion matrix to analyze correct and incorrect classifications incurred by our model, as will be illustrated and elaborated in the following section.
Data Augmentations
In alignment with a majority of FSL works, we stressed the importance of data augmentation in our approach to attain state-of-the-art few-shot classification performances. In light of this, for the training set of the AIDER and CDD datasets, we deployed the following image augmentations: center cropping (by 128 $\times$ 128), resize (by 128 $\times$ 128), random vertical and horizontal flips, random rotations by 90, 180 and 270 deg, random horizontal flips followed by the same aforementioned random rotations, as well as random color jitter, blurring, grayscaling, and random CutMix or Mixup [70]. The probability of all augmentation that are performed in a random manner is set to 0.5. For the training set of the benchmarks, we utilized center cropping and resize (by 32 $\times$ 32 for CIFAR-FS and FC-100 and by 84 $\times$ 84 for miniImageNet and tieredImageNet), random vertical and horizontal flips, random rotations by 90 and 270, grayscale, blur, color jitter and random CutMix or Mixup.
6 Results and Discussions
All implemented algorithms underwent a comprehensive evaluation set of 10 runs, and the results, encapsulated as the Average Accuracy (AA) scores, accompanied by the corresponding standard deviations, were duly recorded. The obtained numerical scores represent the performance of the state-of-the-art (SOTA) methodology and our proposed approach and were established as the foundational benchmarks for the trio of datasets under consideration. In order to provide a thorough and robust evaluation of each SOTA algorithm, barring the triplet network (in the aerial image disaster classification dataset), two distinct evaluation methodologies were employed. These are denoted as 5-way-1 shot and 5-way-5 shot evaluations, providing insight into the performance of the models under varying levels of data sparsity. It is, however, important to note a distinction for the triplet network. Given the unique structure of this network, which compares only a single pair and two pairs of images per episode of training respectively, results are exclusively reported for the 5-way-1 shot evaluation. This restriction aligns with the inherent characteristics of the triplet network, elucidating its capabilities within the stipulated parameters of the experimental setup.
As will be seen in the experimental result section, we will include the results of the ablation study and the image reconstruction of the Bhattacharyya distance component of our ATTBHFA-Net on the benchmarks, while delegating the confusion matrix and GRAD-CAM results on the AIDER and CDD. This is because the latter could be utilized for real-time mission-critical task and require a more thorough assessment of their model’s capability to adapt to seen and novel classes.
6.1 Results on the Benchmark Datasets
[…]
Table 3: The classification (average) accuracy (in $\%$) for the 5-way 1-shot and 5-way 5-shot learning evaluations utilizing our ATTBHFA-Net relative to the SOTAs on the CIFAR-FS. Each method were evaluated for 10 times per setting and the mean were reported. The bolded values denote the best (highest) value obtained, while the underlined values denote the second-best (or second-highest) value obtained.
Methods 5-way-1-shot 5-way-5-shot
[…]
  • *All models are trained with the ResNet12 CNN backbone in a standardized environment.
Table 4: The classification (average) accuracy (in $\%$) for the 5-way 1-shot and 5-way 5-shot learning evaluations utilizing our ATTBHFA-Net relative to the SOTAs on the FC-100. Each method were evaluated for 10 times per setting and the mean were reported. The bolded values denote the best (highest) value obtained, while the underlined values denote the second-best (or second-highest) value obtained.
Methods 5-way-1-shot 5-way-5-shot
[…]
  • *All models, except for AssoAlign which utilized the ResNet18, are trained with the ResNet12 CNN backbone in a standardized environment.
Table 5: The classification (average) accuracy (in $\%$) for the 5-way 1-shot and 5-way 5-shot learning evaluations utilizing our ATTBHFA-Net relative to the SOTAs on the miniImageNet and tieredImageNet. Each method were evaluated for 10 times per setting and the mean were reported. The bolded values denote the best (highest) value obtained, while the underlined values denote the second-best (or second-highest) value obtained.
Datasets
miniImageNet tieredImageNet
Methods 5-way-1-shot 5-way-5-shot 5-way-1-shot 5-way-5-shot
[…]
  • *All models, except for AssoAlign which utilized the ResNet18, are trained with the ResNet12 CNN backbone in a standardized environment.
The main reason why ANROT-HELANet performed better than both the ATTBHFA-Net and the HELA-VFA in the 5-shot setting, despite the ANROT-HELANet utilizing only the Hellinger distance, is the adversarial and natural noise trainings that were involved in the latter. Such a procedure can further enhance the few-shot classification performances, as introducing the adversarial and natural noises in the samples serves as additional data augmentation steps that allows the model to better generalize to unseen classes at a new unprecedented level. However, such effectiveness also requires exposure to as many shot or query samples as possible, similar to the aforementioned observation about 5-shot learning yielding higher metric values in general. Therefore, this explains why ANROT-HELANet in 5-shot performance can achieve higher AA than our proposed approach, but not in 1-shot evaluation. This also sheds light on the possibility of introducing adversarial or natural noise to our model too to enhance its generalization capability in the benchmark setting (but not in the disaster image classification, for reasons already described in the previous section).
6.1.1 Ablation Study
Table 6: Ablation study results on the ATTBHFA-Net for the various configurations tabulated below on the miniImageNet dataset. Bold values denote the optimal values obtained out of all the configurations attempted.
Configuration 5-way-1-shot 5-way-5-shot
[…]
We conducted an ablation study investigating the role of each of our ATTBHFA-Net’s novel component on 1-shot and 5-shot performance on the miniImageNet dataset. The tabulated results are displayed in Table 6. For a more holistic evaluation, we also compared our model’s performances when we replaced the $\ell_{BHS}$ with the individual Hellinger similarity softmax loss $\ell_{Hesim}$ and the KL divergence $\ell_{KL}$, which is the most widely utilized variational loss function and probability distribution comparison metric. Take note that throughout the tables, the configuration labeled “W/o $\ell_{BHS}$ or $\ell_{KL}$ or $\ell_{Hesim}$” and “W/o Att, $\ell_{BHS}$ or $\ell_{KL}$ or $\ell_{Hesim}$” denote ATTBHFA-Net training using only the cross-entropy loss. Except for the “Attention + $\ell_{BHS}$” and “No Attention + $\ell_{BHS}$” configuration, the value of the ablation study are obtained from [40], as the ablation study conducted here are almost the same as that in the latter’s work.
We can observe from the tables that relatively lower values are obtained for all shot settings when the attention mechanism was not included, regardless of whether $\ell_{KL}$, $\ell_{Hesim}$ or $\ell_{BHS}$ is used (or when both are not used). This clearly emphasizes the role of attention in focusing only on relevant regions and key features within the respective classes of the images to enhance the model’s classification. We also noted that regardless of whether or not attention is included, the classification values obtained using $\ell_{KL}$ are generally lower than those of $\ell_{Hesim}$ and $\ell_{BHS}$. One possible reason is that, as already mentioned in [40], the KL divergence is not an actual distance metric, as it does not satisfy the triangle inequality. Consequently, for a pair of not-so-distant distribution configurations in which the values are closely similar to each other (e.g., $(p,q)$, $(p^{\prime},q^{\prime})$ where $p=q^{\prime}$ and $q=p^{\prime}$), $\ell_{KL}$ might return different values and subsequently affects the network’s feature clustering capability. Another reason is that the loss functions $\ell_{Hesim}$ or $\ell_{BHS}$ are closely related as the latter only differs from the former by an additional $0.5\{v_{(Q,i)},c_{(Q,j)}\}^{2}$ term, which encapsulates the additional Bhattarcharyya coefficient contribution to $\ell_{BHS}$. Since the configurations with $\ell_{Hesim}$ were shown to outperform those with $\ell_{KL}$, the aforementioned value trends naturally follow.
6.1.2 Image Reconstruction Analysis
Figure 7: The miniImageNet reconstructed images obtained from the Vanilla VAE architectures using the KL divergence (Top Left), Wasserstein distance (Top Right), Hellinger distance (Bottom Left), and the ATTBHFA-Net Bhattarcharyya distance component (Bottom Right) (zoom in for the best view).
We also conducted a study evaluating the quality of the reconstructed miniImageNet images using the Bhattacharyya distance component of our generative model relative to the selected existing VAEs. Some instances of the latter include the original vanilla VAE (based on the KL divergence) [47], the vanilla Wasserstein VAE (WAE) ([31],[88]), and the Hellinger VAEs ([40], [41]). In our analysis, we only utilized the test set of the miniImageNet, as not only meta-learning was not involved in the reconstruction process, but such study is for comparing the effect of employing different distance metrics on the ELBO optimization in the quality of the reconstructed images.
The metric utilized for the comparison between the reconstructed and the original images is the Frechet Inception Distance (FID) [89], which computes the distances between the feature vectors of the two image sets using the InceptionV3 model [90]. A lower FID score implies that two image sets have more similar statistics, with a score of 0 denoting a perfect identical match between the image sets. The first 10000 images from the miniImageNet meta-test dataset were utilized for the image reconstruction. The FID is mathematically described by
[…]
where $(\mu_{r},\Sigma_{r})$ and $(\mu_{g},\Sigma_{g})$ are the means and covariances of the real and generated images respectively, and $Tr$ denote the trace of the matrix in the parenthesis.
Fig.7 depicts some of the selected reconstructed miniImageNet images using the original VAE (top left), the WAE (top right), the Hellinger VAE (bottom left), and the Bhattarcharyya distance component of our ATTBHFA-Net (bottom right) respectively. At the same time, Table 7 represents the FID values obtained for the respective method. As we can see by zooming into the bottom left of Fig.7, the HELA-VFA visually outputted higher quality reconstructed images than the other two VAE models (KL-divergence and Wasserstein distance). In particular, the images that illustrated a sharp contrast is of the tank (row 4 column 2), the concert player (row 5 column 4), a group of people walking (row 4 column 6 and row 6 column 6), and lastly the singer (row 6 column 3). Still, the Bhattarcharyya distance (bottom right of Fig.7) manages to bring the image quality slightly to a new higher-level, having a FID value of 2.49 and corresponding to a difference of 0.26 relative to the HELA-VFA.
The above observations can be partly attributed to the Bhattarcharyya distance preserving richer latent representations and attaining better balancing between overlap and separability of the feature distributions than the other f-divergences. For the former, this is due to the mathematical definition of the Bhattarcharyya coefficient which (like the Hellinger distance) serves to optimize any overlap (no matter how small) between feature distribution and hence establishing a more structured and regularized latent space with a smoother manifold. By sampling from such space or manifold, higher-quality and coherent images can be reconstructed, leading to a lower FID score. For the latter, we note again that the Bhattacharyya distance (like the Hellinger distance) doesn’t suffer from the instability issue inherent in KL divergence. This means that the encoder could ignore the input and just output the prior features to mitigate the infinities that could occur when the model penalizes any latent distribution that is considered highly unlikely, leading to poor reconstruction of the generated images. Although the Hellinger distance and the Bhattarcharyya distance are closely related, as shown in section 4.1.4, the Hellinger distance linearly decreases as the overlap between distributions increases, while the Bhattarcharyya distance has logarithmically amplification even under small overlap of the distribution. This means that for dissimilar latent feature representation, more discriminative gradient signals could be given than that of the Hellinger distance, enabling better contrastive aggregation of the class feature distributions and subsequently leading to sharper reconstruction of the images.
Table 7: The FID values obtained from the selected VAE architectures for the miniImageNet dataset. Take note that the quality of the reconstructed images improves as the values decrease.
VAE Model Distance utilized FID
Vanilla VAE [47] KL Divergence 3.43
WAE ([31], [88]) Wasserstein Distance 3.38
HELA-VFA ([40]) Hellinger Distance 2.75
ATTBHFA-Net (Bhattarcharyya component) Bhattarcharyya Distance 2.49
6.2 Results on the Aerial Disaster Image Classification Datasets
Table 8: The classification (average) accuracy (in $\%$) using the 5-way 1-shot and 5-way 5-shot learning evaluation on our AIDER subset simulation. Each method were evaluated for 10 times per setting and the mean were reported. The bolded values denote the best (highest) value obtained, while the underlined values denote the second-best (or second-highest) value obtained.
Methods 5-way-1-shot 5-way-5-shot
[…]
  • *All models, except for RelationNet which utilized U-Net [92], are trained with the ResNet12 CNN backbone in a standardized environment. The triplet network employs the original paper’s specified CNN backbone. The data augmentation techniques as mentioned in the previous section is utilized for all methods.
Table 9: The classification (average) accuracy (in $\%$) using the 5-way 1-shot and 5-way 5-shot learning evaluation on the CDD meta-test dataset. Each method were evaluated for 10 times per setting and the mean were reported. The bolded values denote the best (highest) value obtained, while the underlined values denote the second-best (or second-highest) value obtained.
Methods 5-way-1-shot 5-way-5-shot
[…]
  • *The models were trained under uniform environmental conditions, with the ResNet12 serving as the CNN backbone for all except RelationNet, which employed U-Net, and the triplet network that used the original paper’s CNN backbone.
[…]
Figure 8: Example of a possible class confusion that might arose when utilizing the CDD dataset due to the wider context available. The image on the left and middle is easily confused by our algorithm for fire disaster due to its prominent orange streak and light. The image on the far right depicts an actual fire class for comparison.
Figure 9: Another example of a possible class confusion that might arose when utilizing the CDD dataset due to the wider context available. The image on the left and middle is easily confused by our algorithm for a non-disaster infrastructure class image due to the latter’s old and predominantly weary features, although the infrastructure damage images also feature buildings that does not suffer extensive and obvious visual damge indicators (see middle image). The image on the far right depicts an actual (non-disaster) infrastructure class for comparison.
6.2.1 Confusion Matrices
To elaborate on the previous observations, we analyzed and generated confusion matrices for the two aforementioned datasets, with Fig.10 depicting that for the 5-way-1-shot and 5-way-5-shot evaluation of the AIDER, and Fig.11 depicting the corresponding matrices for the CDD.
Figure 10: The confusion matrix generated for the 5-way-1-shot approach (left) and for the 5-way-5-shot approach (right) for our ATTBHFA-Net on the AIDER. Note that in both the predicted and true label, class 1-5 in the table represents the collapsed building, fire, water disaster, traffic accident, and non-disaster class respectively.
Figure 11: The confusion matrix generated for the 5-way-1-shot approach (left) and for the 5-way-5-shot approach (right) for our ATTBHFA-Net on the CDD. Note that in both the predicted and true label, class 1-5 in the table represents the (non-disaster) forest, earthquake, urban fire, water disaster and drought class respectively.
[…]
Figure 12: Yet another example of a possible class confusion that might arose when utilizing the CDD dataset due to the wider context available. This time the confusion arose due to possible water bodies appearing in image categorized as normal class, as shown on the left of the figure. Consequently, the appearance of such images could deteriorate the effective few-shot classification of the water disaster image class, leading to its low correct predicted accuracy. Since the (non-disaster) infrastructure class is not in the meta-test dataset, the model would assign predicted labels to the class which bears the closest visual features with the latter, such as the drought (some images contained certain amount of water bodies as shown in the third picture of the figure) and the earthquake class (the gravel and concrete slope could be confused from a flooded polluted river), leading to its high confusion rate as depicted in the confusion matrix.
6.2.2 GRAD-CAM
We also highlighted some of the Gradient-weighted Class Activation Mapping (GRAD-CAM) activation map on selected AIDER disaster images Fig. 13 and CDD in Fig.14. Using such maps, we can dive into how our ATTBHFA-Net interprets the respective disaster classes in the respective images in the datasets compared to the selected SOTAs (MatchingNet, RelationNet, and ProtoNet). For the CDD images, we also presented a mix of training class (collapsed building, land disaster) and testing classes (water disaster, fire).
For the AIDER collapsed building image, the key feature that represents the class is the rubble near the orange bulldozer. As we can see in Fig. 13, our ATTBHFA-Net successfully picks up the cue at that region on a broader scale, while the other SOTAs either pick it up on a smaller region (i.e., MatchingNet, RelationNet) or pay attention to the wrong part of the region (i.e., ProtoNet). Likewise, for the fire setting, our ATTBHFA-Net is able to pinpoint the fire hotspot located at the center accurately, whereas for the MatchingNet, the heat map pinpoints the wrong part of the fire image, while the ProtoNet and the RelationNet also pinpoint elevant areas apart from the hotspot. A more challenging scenario emerged from the flood image. As we can observe, the flooded areas are more conspicuous at the bottom of the image. For all of the selected methods, the heatmap is only able to focus on the region partially. The reason may be attributed to the grayish contrast of the flooded features, which can be easily confused for the background houses with gray rooftops.
Figure 13: The grad-cam attention maps for selected FSL approaches on the AIDER images for each of the disaster classes. We excluded the class of non-disaster (normal) images in the analysis.
Figure 14: The grad-cam attention maps for selected FSL approaches on the CDD images for each of the disaster classes. Like AIDER, we excluded the non-disaster (normal) image class in the analysis.
For the CDD GRAD-CAM analysis, we once again found that most of the selected algorithm pinpoints the fire hotspot (to the right of the fire image as in Fig.14) accurately, with MatchingNet and ProtoNet being able to highlight the spot entirely. For ATTBHFA-Net, it seems that the attention maps is more spread out than the two aforementioned methods, while the RelationNet fails to spot the heat region. Our ATTBHFA-Net in this context is able to pinpoint the key water disaster (flood) regions and building rubble regions (infrastructure damages/collapsed building) more conspicuously than the other methods. The key challenge for the algorithm is the detection of key features for the land disaster class). We observed that the key visual cues that distinguished it as the aforementioned class is the displacement of the massive pile of earth, as seen on the bottom of the image. Surprisingly, MatchingNet is able to locate and focus on the region containing the features to a larger extent than the other methods. In fact, our ATTBHFA-Net and, to a smaller extent, the RelationNet, actually highlight the wrong regions which is of the opening in the house.
In summary, the above GRAD-CAM analysis indicated that our ATTBHFA-Net has correctly highlighted the key visual signature for each disaster types in AIDER to a high level of degree, while yielding mixed results for the other two datasets. Although the analysis was done for one image per disaster class for all the datasets, the observations and trends can be extrapolated to the rest of the images, since we have mentioned beforehand that the CDD has more disaster feature variations and contexts than AIDER, which would lead to higher probability of wrong GRAD-CAM highlights regardless of the approaches. This may help to partially explain why ATTBHFA-Net has the highest AA for both the 1-shot and 5-shot approaches in AIDER, and why the overall AA values are lower for the other SOTAs for CDD relative to AIDER.
Lastly, it is also worthwhile to mention that some works such as [93] and [94] has emphasized the importance of utilizing a finer spatial resolution during the classification stages and assessing its impact on the classification performance. The significance of this issue is put forth since in disaster scenarios, the signature features of such disasters typically only occurred within a small region in a bigger captured field of view in the imagery. Although this aspect was not thoroughly explored in this paper as our emphasis was on assessing the algorithmic effectiveness of our few-shot approaches, it is nonetheless an interesting and essential perspective that we will investigate in our future works.
7 Conclusions
Our approach employs the benchmarks and the UAV-based remote sensing images to tackle natural disaster scenarios. We exhibit the efficacy of ATTBHFA-Net in four benchmarks and two different tasks for disaster classification, establishing the latter as a promising benchmark for future investigations in this area. By utilizing an attention mechanism with prototype generation through the Bhattarcharyya-Hellinger metric, we present an inventive method that boosts feature extraction and conserves essential class knowledge pertinent to each task at hand. Our findings make a significant contribution to improving understanding of how cutting-edge techniques can work harmoniously together to address intricate real-world issues including emergency response situations brought about by natural disasters. In all data sets, our approach in general outperformed the SOTA FSLs in the AA scores, demonstrating the feasibility and effectiveness of our FSL design in allowing better classification of a wide variety of image-based disaster scenery. In our future work, the proposed ATTBHFA-Net architecture will be tested in real-time disaster identification scenario (e.g., processing images of aftermath of flooding) for validation purposes, and alternative methods will be explored to further improve the model performance, which includes, for example, such GAN-based augmentation [95, 96].
Appendix A Derivation of the Modified ELBO ($ELBO^{*}$) for Bhattarcharyya Distance (and Coefficient)
In this appendix section, we provide a derivation for the modified ELBO ($ELBO^{*}$) that relates to the Bhattarcharyya distance $D_{B}$ (and hence the coefficient $BC$). We start with the definition of $BC$ as
$BC=\int\sqrt{q_{\phi}(z|\mathcal{S})p_{\theta}(z|\mathcal{T})}dz$ (21)
The derivation works best if we consider the Bhattarcharryya distance $D_{B}$ directly, which is defined as the negative of the natural logarithm of $BC$ as $D_{B}=-ln(BC)$. From this we take the derivative with respect to $z$ on both sides to get
$\frac{d}{dz}\left(BC\right)=\sqrt{q_{\phi}(z|\mathcal{S})p_{\theta}(z|\mathcal{T})}.$ (22)
Taking natural logarithm on both sides and expanding out the right-hand side, we get
$ln\left(\frac{d}{dz}\left(e^{-D_{B}}\right)\right)=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T})).$ (23)
We now use the property that $\int q_{\phi}(z|\mathcal{S})dz=1$ and introduced it on the left-hand side of the equation to get
$ln\left(\frac{d}{dz}\left(e^{-D_{B}}\int q_{\phi}(z|\mathcal{S})dz\right)\right)=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T})),$ (24)
where in general, $D_{B}$ is dependent on z, but in the VAE context, the posterior and the prior distributions have means and variances independent of $z$, so we can bring the $e^{-D_{B}}$ term in the integral to obtain
$ln\left(\frac{d}{dz}\int q_{\phi}(z|\mathcal{S})e^{-D_{B}}dz\right)=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T})),$ (25)
$ln\left(q_{\phi}(z|\mathcal{S})e^{-D_{B}}\right)=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T})).$ (26)
Expanding out the left-hand side, taking note that $ln(e^{-D_{B}})=-D_{B}$,
$ln\left(q_{\phi}(z|\mathcal{S})\right)-D_{B}=\frac{1}{2}ln(q_{\phi}(z|\mathcal{S}))+\frac{1}{2}ln(p_{\theta}(z|\mathcal{T}))$ (27)
and multiplying both sides by $q_{\phi}(z|\mathcal{S})$, followed by integrating both sides with respect to dz, we get
$\begin{split}\int q_{\phi}(z|\mathcal{S})ln\left(q_{\phi}(z|\mathcal{S})\right)dz-\int D_{B}q_{\phi}(z|\mathcal{S})dz=\\ \frac{1}{2}\int q_{\phi}(z|\mathcal{S})ln(q_{\phi}(z|\mathcal{S}))dz+\frac{1}{2}\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz.\end{split}$ (28)
We note that $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz$ is nothing but the reconstruction term. Rearranging,
$\frac{1}{2}\int q_{\phi}(z|\mathcal{S})ln\left(q_{\phi}(z|\mathcal{S})\right)dz-\frac{1}{2}\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz=\int D_{B}q_{\phi}(z|\mathcal{S})dz,$ (29)
$\int q_{\phi}(z|\mathcal{S})ln\left(q_{\phi}(z|\mathcal{S})\right)dz-\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz=2\int D_{B}q_{\phi}(z|\mathcal{S})dz.$ (30)
Inserting $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ on the left-hand side while shifting the reconstruction term to the right,
$\begin{split}\int q_{\phi}(z|\mathcal{S})ln\left(q_{\phi}(z|\mathcal{S})\right)dz+\left(\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz\right)=\\ \int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz+2\int D_{B}q_{\phi}(z|\mathcal{S})dz,\end{split}$ (31)
and we can obtain the negative of the ELBO as $ELBO=\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz=-\int q_{\phi}(z|\mathcal{S})ln\left(\frac{p_{\theta}(z|\mathcal{S},\mathcal{T})}{q_{\phi}(z|\mathcal{S})}\right)dz$,
$\begin{split}\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-ELBO=\\ \int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz+2\int D_{B}q_{\phi}(z|\mathcal{S})dz.\end{split}$ (32)
We make use of a property that is useful (especially during the image reconstruction process): Since $D_{B}$ and thus $BC$ is measured via comparing the distribution $p_{\theta}(z|\mathcal{T})$ and $q_{\phi}(z|\mathcal{S})$, and in VAE the latter can be expressed in terms of the means and covariances of the distribution, all independent of $z$ as they depend on the original input images, we can bring $D_{B}$ out of the integral and utilize $\int q_{\phi}(z|\mathcal{S})=1$ again to get
$\begin{split}\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-ELBO=\\ \int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz+2D_{B}.\end{split}$ (33)
Finally, define the $ELBO^{*}$ as $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-ELBO$, we arrived at
$\begin{split}ELBO^{*}=\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{T}))dz+2D_{B}=\int q_{\phi}(z|\mathcal{S})\textrm{ln}\left(p_{\theta}(z|\mathcal{T})\right)dz-2ln(BC)\end{split}$ (34)
which is equation 6.
Appendix B Another Derivation of the Modified ELBO ($ELBO^{\prime\prime}$) for the Hellinger Distance
In this appendix section, we provide another derivation for the modified ELBO ($ELBO^{\prime\prime}$) that relates to the square of the Hellinger distance $D_{H}^{2}$ (which has already been in [40] and [41]). We start with the definition of the $D_{H}^{2}$ as
$D_{H}^{2}=1-\int\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz.$ (35)
Switching the terms on each side to get
$\int\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz=1-D_{H}^{2}.$ (36)
Take derivative with respect to $z$ on both sides,
$\frac{d}{dz}\int\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}dz=\frac{d}{dz}(1-D_{H}^{2}),$ (37)
$\sqrt{p_{\theta}(z|\mathcal{T})q_{\phi}(z|\mathcal{S})}=\frac{d}{dz}(1-D_{H}^{2}),$ (38)
and taking logarithm on both sides,
$\frac{1}{2}log(p_{\theta}(z|\mathcal{T}))+\frac{1}{2}log(q_{\phi}(z|\mathcal{S}))=log\left(\frac{d}{dz}(1-D_{H}^{2})\right),$ (39)
$log(p_{\theta}(z|\mathcal{T}))+log(q_{\phi}(z|\mathcal{S}))=2log\left(\frac{d}{dz}(1-D_{H}^{2})\right).$ (40)
Now we apply an analogous argument as in the previous derivation: Since $D^{2}_{H}$ is measured via comparing the distribution $p_{\theta}(z|\mathcal{T})$ and $q_{\phi}(z|\mathcal{S})$, and in VAE the latter can be expressed in terms of the means and covariances of the distribution, all independent of $z$ as they depend on the original input images, $log\left(\frac{d}{dz}(1-D_{H}^{2})\right)$ is independent of $z$ and we can insert $\int q_{\phi}(z|\mathcal{S})dz=1$ in the parenthesis of $(1-D_{H}^{2})$ to get
$\begin{split}log(p_{\theta}(z|\mathcal{T}))+log(q_{\phi}(z|\mathcal{S}))=2log\left(\frac{d}{dz}(1-D_{H}^{2})\int q_{\phi}(z|\mathcal{S})dz\right)\\ =2log\left(\frac{d}{dz}\int(1-D_{H}^{2})q_{\phi}(z|\mathcal{S})dz\right)\\ =2log\left((1-D_{H}^{2})q_{\phi}(z|\mathcal{S})\right)\\ =2log(1-D_{H}^{2})+2log\left(q_{\phi}(z|\mathcal{S}\right).\end{split}$ (41)
Bringing $2log(q_{\phi}(z|\mathcal{S})$ to the left-hand side of the equation,
$\begin{split}log(p_{\theta}(z|\mathcal{T}))-log(q_{\phi}(z|\mathcal{S}))=2log(1-D_{H}^{2}).\end{split}$ (42)
Multiplying both sides by $q_{\phi}(z|\mathcal{S})$ and integrate them with respect to $z$,
$\begin{split}\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})log(q_{\phi}(z|\mathcal{S}))dz=2\int q_{\phi}(z|\mathcal{S})log(1-D_{H}^{2})dz.\end{split}$ (43)
Use the property that $log(1-D_{H}^{2})$ is independent of $z$ to bring it out of the integral on the right:
$\begin{split}\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})log(q_{\phi}(z|\mathcal{S}))dz=2\left(\int q_{\phi}(z|\mathcal{S})dz\right)log(1-D_{H}^{2})\\ =2log(1-D_{H}^{2}),\end{split}$ (44)
where in the last line we made use of $\int q_{\phi}(z|\mathcal{S})dz=1$ again. Recognizing that $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz$ is the reconstruction term, we insert $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ on the left-hand side while shifting the reconstruction term to the right,
$\begin{split}\left(\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz\right)-\int q_{\phi}(z|\mathcal{S})log(q_{\phi}(z|\mathcal{S}))dz=\\ -\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+2log(1-D_{H}^{2}),\end{split}$ (45)
$\begin{split}\left(\int q_{\phi}(z|\mathcal{S})log\left(\frac{p_{\theta}(z|\mathcal{S},\mathcal{T})}{q_{\phi}(z|\mathcal{S})}\right)dz-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz\right)=\\ -\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+2log(1-D_{H}^{2}),\end{split}$ (46)
$\begin{split}ELBO-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz=-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+2log(1-D_{H}^{2}),\end{split}$ (47)
and we finally arrived at
$\begin{split}ELBO^{\prime\prime}=-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{T}))dz+log(1-D_{H}^{2})^{2},\end{split}$ (48)
which is equation 9. The modified $ELBO^{\prime\prime}$ reads $ELBO-\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$. Some remarks are in order:
  • The reconstruction term in the relationship between $ELBO^{*}$ and $D_{B}$ involved a natural logarithm of $p_{\theta}(z|\mathcal{T})$, while the reconstruction term in the relationship between $ELBO^{\prime\prime}$ and $D^{2}_{H}$ involved a logarithm of base 10 of $p_{\theta}(z|\mathcal{T})$. This is because in our derivation of the relationship between $ELBO^{*}$ and $D_{B}$, we encountered a term of the form $ln(e^{-D_{B}})$, which if expressed in log base 10 could not be reduced to $D_{B}$.
  • The reconstruction term in the relationship between $ELBO^{*}$ and $D_{B}$ is positive on the right-hand side of the equation while the reconstruction term in the relationship between $ELBO^{\prime\prime}$ and $log(1-D_{H}^{2})^{2}$ is negative on the right-hand side of the equation.
  • The modified ELBO in equation 6 ($ELBO^{*}$) involved subtraction by the ELBO term while the modified ELBO in equation 9 $($$ELBO^{\prime\prime}$) involved subtraction directly from ELBO. This means that in the overall optimization objective as shown in equation 11, the original ELBO term cancels out and the modified ELBO summation in the maximization of the objective mainly involved the reconstruction term $\int q_{\phi}(z|\mathcal{S})log(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$ and $\int q_{\phi}(z|\mathcal{S})ln(p_{\theta}(z|\mathcal{S},\mathcal{T}))dz$. This also implies that the summation involved the term containing the square of the Hellinger distance $log(1-D_{H}^{2})^{2}$ getting subtracted by twice of the Bhattarcharyya coefficient $2ln(BC)$.
[…]
Acknowledgment
This research/project is supported by the Civil Aviation Authority of Singapore and NTU under their collaboration in the Air Traffic Management Research Institute. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Civil Aviation Authority of Singapore.
Declarations
  • Funding Information: This research is funded under the ATM-CAAS Leaders’ Track Scholarship, under the NTU ATMRI-CAAS collaboration.
  • Competing interests: On behalf of all authors, the corresponding author states that there is no conflict of interest.
  • Ethical Approval for Human and /or Animals Research: Not applicable.
  • Availability of Data and Materials: All data will be made available upon request.
  • Author contributions (CreDiT Authorship Statement): Gao Yu Lee: Conceptualization, Investigation, Methodology, Software, Writing - original draft, Writing- review &\& editing. Tanmoy Dam: Conceptualization, Investigation, Supervision, Writing- review &\& editing. Md Mefahul Ferdaus: Supervision, Writing- review &\& editing. Daniel Puiu Poenar: Supervision. Vu N. Duong: Supervision.